OpenStudy (elcna):

Help please? The epression (3*x)*5 is evaluated at x=-1/5, the associative property of multiplication is applied, and one multiplication is done. Which is the result? A. (-3/5)*5 B. 3*(-1) C. 5*(-3/5) D. -3*(1)

OpenStudy (anonymous):

what do those asterisks mean?

OpenStudy (anonymous):

is it multiplication ?

OpenStudy (elcna):

Yes, the stars mean multiplication @satellite73

OpenStudy (anonymous):

oh ok if you replace \(x\) by \(-\frac{1}{5}\) you get \[3\times \left(-\frac{1}{5}\right)\times 5\]

OpenStudy (anonymous):

if you multiply the last two terms first, this is the same as \[3\times (-1)\]

OpenStudy (elcna):

Thank you!! @satellite73

OpenStudy (anonymous):

yw